### v1

@@ 3. The $\ell^{p}$-Bott-Dirac construction and twisted ghostly ideals
To distinguish, we denote the geometric ideal and the ghostly ideal of $C^{*}(X,\mathcal{H}_{X,E})$ as $\mathcal{I}_{E}(X,U)$ and $\mathcal{G}_{E}(X,E)$, respectively. By 3.11, we can then restrict the index map on the geometric ideal and ghostly ideal. Thus, we have the following commuting diagram
@@
Analogous to 3.8, we also have the following proposition for ideals. We should mention that the proof of 3.12 can also be applied to 3.8.
@@
We shall only prove the $K_{0}$-case, and the $K_{1}$-case follows from the suspension argument. The proof for $\mathcal{I}(X,U)$ also applies for $\mathcal{G}(X,U)$, thus we shall focus on $\mathcal{I}(X,U)$.
@@
Notice that for each $k\in\mathbb{N}\cup\{\infty\}$, $F^{(k)}=(F^{(k)}_{n})_{n\in\mathbb{N}}$ is in the multiplier of $\mathcal{I}_{E}(X,U)$ by 3.3. Then for any projection $P\in\mathcal{I}(X,U)$, we can define an index map
@@
For each $P]\in K_{0}(\mathcal{I}(X,U))$ [^1] By [HIT2020, Corollary 2.7.4], a $K_{0}$-theory element of $\mathcal{I}(X,U)$ can be represented by a single projection $[P]$., we write $\textup{Ind}^{(k)}([P])=[(p^{(k)}_{0},\cdots,p^{(k)}_{n},\cdots)]-[q_{0},\cdots,q_{n}\cdots]$ for each $k\in\mathbb{N}\cup\{\infty\}$, where $q_{n}$ does not depend on $k$ but only depends on $[P]$ by definition. For simplicity of notation, we write
@@
for all $n\in\mathbb{N}$. Thus, $t\mapsto F_{n}^{(k+t)}$ is equi-continuous as $n$ varies in $\mathbb{N}$. We can built a path
@@
where $p_{k}$ is the one-dimensional projection onto the kernel of $B$. We denote $P]=[(P_{0},\cdots,P_{n},\cdots)]$ and $\textup{Ind}^{(\infty)}([P])=[(Q_{0},\cdots,Q_{n},\cdots)]$. By using the homotopy $(P_{E_{n}}(t))_{t\in[0,1]}$, one can check that $Q_{n}$ is homotopic to $P\otimes p_{k}$. However, the homotopy is not equi-continuous as the dimension of $E_{n}$ increases. The reason comes from that we replaced the parameter $s$ by $s+s_{n}$ for each $E_{n}$. One can check that $\|F_{E_{n}}(t)-F_{E_{n}}(t^{\prime})\|\leq|f(\frac{2}{t(s+s_{n})})-f(\frac{2}{t^{\prime}(s+s_{n})})|\leq(s+s_{n})|t-t^{\prime}|$ the Lipschitz constant is depended on $n$.
However, we can use a ”stacking argument” as in [HIT2020, Proposition 12.6.3] to get around this problem. Define $0,1]\mapsto\mathcal{I}_{E}(X,U)$ by $t\mapsto P(t)=(P_{0}(t),\cdots,P_{n}(t),\cdots)$ where $P_{n}(t)$ is the homotopy between $Q$ and $P\otimes p_{k}$ for each $n\in\mathbb{N}$. Let $N_{n}=[s+2s_{n}]+1$. For each $t\in[0,1]$, we define $P^{N}(t)=(P^{N}_{n}(t))\in\mathcal{I}_{E}(X,U)$
@@
Since the K-theory of $A_{\mathcal{I}(U)}(X,E)$ and $A_{\mathcal{G}(U)}(X,E)$ only depends on their asymptotic behavior as $s\to\infty$, [HIT2020, Lemma 6.4.11], and 3.13 shows that these two algebras have the same asymptotic behavior, the conclusion thus follows naturally. We leave the details to the reader. ∎
@@
The theorem follows directly from the diagram (2), 3.12, 3.13 and a diagram chasing argument. ∎
@@
For $p\in[1,\infty)$. Let $L^{p}=L^{p}[0,1]$ denote the $L^{p}$-space on $([0,1],m)$ equipped with the canonical Lebesgue measure $m$. By [Banachspace, Proposition III.A.1], any separable $L^{p}$-space $L^{p}(\Omega,\mu)$ is isometric to the $p$-direct sum of $L^{p}$ and $\ell^{p}$. Thus, $L^{p}[0,1]$ serves as one of the most representative $L^{p}$-spaces.
@@
Suppose that $X=\bigsqcup_{n}X_{n}$ is a sparse space that admits a uniformly coarse embedding $\{f_{n}:X_{n}\to L^{p}\}$. Since $W$ is dense in $L^{p}$, without loss of generality, we may assume that $X$ coarsely embeds into $W=\bigcup_{n\in\mathbb{N}}W_{n}$. Since $X_{n}$ is a finite set, its image $f_{n}(X_{n})$ is a finite subset of $W$, there exists a sufficiently large integer $k_{n}$ such that $f_{n}(X_{n})\subseteq W_{k_{n}}$. Identifying $W_{k_{n}}$ with the subspace of $\ell^{p}$ supported on the first $2^{k_{n}}$ coordinates, we see that $f_{n}$ effectively embeds $X_{n}$ into $\ell^{p}$. Thus, the existence of a coarse embedding into $L^{p}$ implies the existence of a coarse embedding into $\ell^{p}$. By [Banachspace, Proposition III.A.1], any separable $L^{p}(\Omega,\mu)$ is isometric to the direct sum of $L^{p}[0,1]$ and $\ell^{p}$. Thus, our main result holds if we replace $\ell^{p}$ by any general separable $L^{p}$-spaces.
@@
The conclusion of Theorem 2.6 holds true if the assumption of coarse embeddability into $\ell^{p}$ is replaced by coarse embeddability into a $L^{p}$-space ($1\leq p<\infty$).∎
@@
We should mention that 3.16 can also be proved by the finite representability of $L^{p}$ in $\ell^{p}$. see [Banachspace, Section III.E.15]. Recall that a Banach space $X$ is finitely representable in $Y$ if any finite-dimensional subspace of $X$ is $(1+\varepsilon)$-isomorphic to a subspace of $Y$ for any $\varepsilon>0$.
As a consequence of [Banachspace, Section III.E.15], the family of all finite-dimensional subspaces of $L^{p}$ admits a uniformly coarse embedding into $\ell^{p}$. Thus, a sparse space $X$ admits a coarse embedding into $L^{p}$ implies that $X$ admits a coarse embedding into $\ell^{p}$.
@@ 4.1. Applications
In [GWZ2025], a relative version of the coarse Baum-Connes conjecture was introduced. In fact, our main result has the following corollary:
@@
For $d\geq 0$, we define $P_{d}(X)$ as the Rips complex of $X$ at scale $d$. Its metric is defined as the hemispherical metric given in [HIT2020, Definition 7.2.8]. Let $C^{*}_{L}(P_{d}(X))$ be the localization algebra associated with $P_{d}(X)$. Define $C^{*}_{L,Y}(P_{d}(X))$ as the subalgebra of $C^{*}_{L}(P_{d}(X))$ consisting of all functions taking values in $\mathcal{G}(P_{d}(X),U_{Y})$. This is naturally an ideal. We consider the quotient
@@
By 2.8, the right-hand side is a coarse equivalence invariant, and thus is isomorphic to $K_{*}(C^{*}_{Y,\infty}(X))$. The relative assembly map for $(X,Y)$ is defined as the inductive limit of these index maps as $d\to\infty$:
@@
If the subspace $Y$ is chosen to be a bounded set, then the relative coarse Baum-Connes conjecture is called the boundary coarse Baum-Connes conjecture, see .
@@
where $A_{Y,L,d}=C^{*}_{L,Y}(P_{d}(X))$, $A_{L,d}=C^{*}_{L}(P_{d}(X))$, $A_{L,Y,\infty,d}=C^{*}_{L,Y,\infty}(P_{d}(X))$, $A_{Y}=\mathcal{G}(X,U_{Y})$, $A=C^{*}(X)$, and $A_{Y,\infty}=C^{*}_{Y,\infty}(X)$. Since $X$ admits a coarse embedding into $\ell^{p}$, by [WXYZ2024], we have that the maps (2) and (5) are isomorphism for both $i=0,1$. By 2.6, $K_{*}(\mathcal{G}(X,U_{Y}))\cong K_{*}(\mathcal{I}(X,U_{Y}))\cong K_{*}(C^{*}(Y))$. Thus, the maps (1) and (4) are identified with the coarse Baum-Connes assembly map for $Y$, which are also isomorphisms since $Y$ also admits a coarse embedding into $\ell^{p}$. Then the corollary follows from the Five Lemma. ∎
The following definition is introduced by M. Braga, I. Farah, and A. Vignati in [BFV2024].
Definition 4.2 ([BFV2024]).
@@
    If we replace $\mathcal{P}_{Fin}(\prod^{u}_{\mathcal{I}}C^{*}_{u}(X))$ by $\prod^{u}_{\mathcal{I}}C^{*}_{u}(X)$, then we shall say $X$ has operator norm localization, (abbrev. ONL, see [CTWY2008]).
We also have the following corollary.
@@
To prove 4.3, we need the following lemma.
@@
Let $X$ be a sparse metric space. All ghost projections in $C^{*}(X)$ are compact if and only if the canonical inclusion $i:\mathcal{K}\to\mathcal{G}$ induces an isomorphism $i_{*}:K_{0}(\mathcal{K})\to K_{0}(\mathcal{G})$, where $\mathcal{G}=\mathcal{G}(X,X)$ is the ghostly ideal associated with the invariant set $X\subseteq\beta X$.
@@
The $(\Rightarrow)$ part. For any projection $P\in M_{n}(\mathcal{G})$, since the Roe algebra is quasi-stable, there exists a projection $P^{\prime}\in\mathcal{G}$ such that $P]=[P^{\prime}]\in K_{0}(\mathcal{G})$. By assumption, $P^{\prime}$ is compact, thus $P^{\prime}]\in K_{0}(\mathcal{K})$ and $i_{*}([P^{\prime}])=[P]$. Thus $i_{*}$ is surjective. Since $X$ is sparse, the map $i_{*}$ is injective by [OOY2009, Proposition 2.10 & Remark 2.12]. This shows that $i_{*}$ is an isomorphism.
The $(\Leftarrow)$ part. Assume for a contradicition that $P\in C^{*}(X)$ is a non-compact ghost projection. Say $X=\bigsqcup_{n\in\mathbb{N}}X_{n}$ and $N_{n}=\#X_{n}$. Notice that $\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})$ is a subalgebra of $C^{*}(X)$ and $\mathcal{K}(\mathcal{H}_{X})\cap\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})=\bigoplus_{n\in\mathbb{N}}M_{N_{n}}(\mathcal{K})$. Moreover, the quotient algebras are canonically isomorphic to each other, i.e.,
@@
Assume for a contradiction that $X$ does not have $ONL_{\mathcal{P}_{Fin}}$. By [BFV2024, Theorem 1.3], there exists a sparse subspace $X^{\prime}\subseteq X$ such that there exists a non-compact ghost projection in $C^{*}_{u}(X^{\prime})$. Without loss of generality, we may assume that $X$ itself is sparse and and $P\in C^{*}_{u}(X)$ is a non-compact ghost projection. Take $p\in\mathcal{K}(\mathcal{H})$ to be a rank one-projection. Then the tensor product $P\otimes p$ defines a non-compact ghost projection in $C^{*}(X)$. By 4.4, the map $i_{*}$ is not an isomorphism on $K_{0}$. This leads to a contradiction to 2.6. ∎
We should mention that 4.3 provides a large class of examples satisfying the property $ONL_{\mathcal{P}{Fin}}$, including all metric spaces that admit a coarse embedding into Hilbert space.

### v2

@@ 3. The $\ell^{p}$-Bott-Dirac construction and twisted ghostly ideals
To distinguish, we denote the geometric ideal and the ghostly ideal of $C^{*}(X,\mathcal{H}_{X,E})$ as $\mathcal{I}_{E}(X,U)$ and $\mathcal{G}_{E}(X,E)$, respectively. By 3.12, we can then restrict the index map on the geometric ideal and ghostly ideal. Thus, we have the following commuting diagram
@@
Analogous to 3.9, we also have the following proposition for ideals. We should mention that the proof of 3.13 can also be applied to 3.9.
@@
We shall only prove the $K_{0}$-case, and the $K_{1}$-case follows from the suspension argument. The proof for $\mathcal{I}(X,U)$ also applies to $\mathcal{G}(X,U)$, thus we shall focus on $\mathcal{I}(X,U)$.
@@
Notice that for each $k\in\mathbb{N}\cup\{\infty\}$, $F^{(k)}=(F^{(k)}_{n})_{n\in\mathbb{N}}$ is in the multiplier of $\mathcal{I}_{E}(X,U)$ by 3.4. Then for any projection $P\in\mathcal{I}(X,U)$, we can define an index map
@@
For each $P]\in K_{0}(\mathcal{I}(X,U))$ [^2] By [Willett and Yu, 2020, Corollary 2.7.4], a $K_{0}$-theory element of $\mathcal{I}(X,U)$ can be represented by a single projection $[P]$., we write $\textup{Ind}^{(k)}([P])=[(p^{(k)}_{0},\cdots,p^{(k)}_{n},\cdots)]-[q_{0},\cdots,q_{n}\cdots]$ for each $k\in\mathbb{N}\cup\{\infty\}$, where $q_{n}$ does not depend on $k$ but only depends on $[P]$ by definition. For simplicity of notation, we write
@@
for all $n\in\mathbb{N}$. Thus, $t\mapsto F_{n}^{(k+t)}$ is equi-continuous as $n$ varies in $\mathbb{N}$. We can build a path
@@
where $p_{k}$ is the one-dimensional projection onto the kernel of $B$. We denote $P]=[(P_{0},\cdots,P_{n},\cdots)]$ and $\textup{Ind}^{(\infty)}([P])=[(Q_{0},\cdots,Q_{n},\cdots)]$. By using the homotopy $(P_{E_{n}}(t))_{t\in[0,1]}$, one can check that $Q_{n}$ is homotopic to $P\otimes p_{k}$. However, the homotopy is not equi-continuous as the dimension of $E_{n}$ increases. The reason comes from the fact that we replaced the parameter $s$ by $s+s_{n}$ for each $E_{n}$. One can check that $\|F_{E_{n}}(t)-F_{E_{n}}(t^{\prime})\|\leq|f(\frac{2}{t(s+s_{n})})-f(\frac{2}{t^{\prime}(s+s_{n})})|\leq(s+s_{n})|t-t^{\prime}|$ the Lipschitz constant is depended on $n$.
However, we can use a ”stacking argument” as in [Willett and Yu, 2020, Proposition 12.6.3] to get around this problem. Define $0,1]\mapsto\mathcal{I}_{E}(X,U)$ by $t\mapsto P(t)=(P_{0}(t),\cdots,P_{n}(t),\cdots)$ where $P_{n}(t)$ is the homotopy between $Q$ and $P\otimes p_{k}$ for each $n\in\mathbb{N}$. Let $N_{n}=[s+2s_{n}]+1$. For each $t\in[0,1]$, we define $P^{N}(t)=(P^{N}_{n}(t))\in\mathcal{I}_{E}(X,U)$
@@
Since the K-theory of $A_{\mathcal{I}(U)}(X,E)$ and $A_{\mathcal{G}(U)}(X,E)$ only depends on their asymptotic behavior as $s\to\infty$, [Willett and Yu, 2020, Lemma 6.4.11], and 3.14 shows that these two algebras have the same asymptotic behavior, the conclusion thus follows naturally. We leave the details to the reader. ∎
@@
The theorem follows directly from the diagram (3), 3.13, 3.15, and a diagram chasing argument. ∎
@@
For $p\in[1,\infty)$. Let $L^{p}=L^{p}[0,1]$ denote the $L^{p}$-space on $([0,1],m)$ equipped with the canonical Lebesgue measure $m$. By [Wojtaszczyk, 1991, Proposition III.A.1], any separable $L^{p}$-space $L^{p}(\Omega,\mu)$ is isometric to the $p$-direct sum of $L^{p}$ and $\ell^{p}$. Thus, $L^{p}[0,1]$ serves as one of the most representative $L^{p}$-spaces.
@@
Suppose that $X=\bigsqcup_{n}X_{n}$ is a sparse space that admits a uniformly coarse embedding $\{f_{n}:X_{n}\to L^{p}\}$. Since $W$ is dense in $L^{p}$, without loss of generality, we may assume that $X$ coarsely embeds into $W=\bigcup_{n\in\mathbb{N}}W_{n}$. Since $X_{n}$ is a finite set, its image $f_{n}(X_{n})$ is a finite subset of $W$, there exists a sufficiently large integer $k_{n}$ such that $f_{n}(X_{n})\subseteq W_{k_{n}}$. Identifying $W_{k_{n}}$ with the subspace of $\ell^{p}$ supported on the first $2^{k_{n}}$ coordinates, we see that $f_{n}$ effectively embeds $X_{n}$ into $\ell^{p}$. Thus, the existence of a coarse embedding into $L^{p}$ implies the existence of a coarse embedding into $\ell^{p}$. By [Wojtaszczyk, 1991, Proposition III.A.1], any separable $L^{p}(\Omega,\mu)$ is isometric to the direct sum of $L^{p}[0,1]$ and $\ell^{p}$. Thus, our main result holds if we replace $\ell^{p}$ by any general separable $L^{p}$-spaces.
@@
The conclusion of 2.6 holds true if the assumption of coarse embeddability into $\ell^{p}$ is replaced by coarse embeddability into an $L^{p}$-space ($1\leq p<\infty$).∎
@@
It is worth noting that the assumption of separability in 3.16 is actually not necessary. We can provide an alternative proof based on the finite representability of general $L^{p}$-spaces in $\ell^{p}$, as discussed in [Wojtaszczyk, 1991, Section III.E.15]. Recall that a Banach space $X$ is finitely representable in $Y$ if for any finite-dimensional subspace $E\subseteq X$ and any $\epsilon>0$, there exists a subspace $F\subseteq Y$ and an isomorphism $T:E\to F$ such that $\|T\|\cdot\|T^{-1}\|<1+\epsilon$.
The essential reason why any $L^{p}$-space (possibly non-separable) is finitely representable in $\ell^{p}$ lies in the approximation by simple functions. Specifically, any function in a finite-dimensional subspace of $L^{p}$ can be approximated by simple functions, which are linear combinations of characteristic functions over disjoint measurable sets. Since the number of basis vectors in the subspace is finite, the approximation involves only finitely many such measurable sets. The characteristic functions on these disjoint sets naturally span a finite-dimensional subspace isometric to $\ell^{p}_{N}$. Consequently, the linear map embedding (from basis to basis) the finite-dimensional subspace of $L^{p}$ into the generated $\ell^{p}_{N}$ is an almost isometry. This implies that the family of all finite-dimensional subspaces of any $L^{p}$-space admits a uniformly coarse embedding into $\ell^{p}$. Thus, coarse embeddability into a general $L^{p}$-space implies coarse embeddability into $\ell^{p}$ for sparse spaces, extending our main result to the non-separable setting.
@@ 4.1. Applications
In Guo et al. [2025], a relative version of the coarse Baum-Connes conjecture was introduced. In fact, our main result has the following corollary:
@@
For $d\geq 0$, we define $P_{d}(X)$ as the Rips complex of $X$ at scale $d$. Its metric is defined as the hemispherical metric given in [Willett and Yu, 2020, Definition 7.2.8]. Let $C^{*}_{L}(P_{d}(X))$ be the localization algebra associated with $P_{d}(X)$. Define $C^{*}_{L,Y}(P_{d}(X))$ as the subalgebra of $C^{*}_{L}(P_{d}(X))$ consisting of all functions taking values in $\mathcal{G}(P_{d}(X),U_{Y})$. This is naturally an ideal. We consider the quotient
@@
By 2.9, the right-hand side is a coarse equivalence invariant, and thus is isomorphic to $K_{*}(C^{*}_{Y,\infty}(X))$. The relative assembly map for $(X,Y)$ is defined as the inductive limit of these index maps as $d\to\infty$:
@@
If the subspace $Y$ is chosen to be a bounded set, then the relative coarse Baum-Connes conjecture is called the boundary coarse Baum-Connes conjecture, see Finn-Sell and Wright [2014].
@@
where $A_{Y,L,d}=C^{*}_{L,Y}(P_{d}(X))$, $A_{L,d}=C^{*}_{L}(P_{d}(X))$, $A_{L,Y,\infty,d}=C^{*}_{L,Y,\infty}(P_{d}(X))$, $A_{Y}=\mathcal{G}(X,U_{Y})$, $A=C^{*}(X)$, and $A_{Y,\infty}=C^{*}_{Y,\infty}(X)$. Since $X$ admits a coarse embedding into $\ell^{p}$, by Wang et al. [2024], we have that the maps (2) and (5) are isomorphisms for both $i=0,1$. By 2.6, $K_{*}(\mathcal{G}(X,U_{Y}))\cong K_{*}(\mathcal{I}(X,U_{Y}))\cong K_{*}(C^{*}(Y))$. Thus, the maps (1) and (4) are identified with the coarse Baum-Connes assembly map for $Y$, which are also isomorphisms since $Y$ also admits a coarse embedding into $\ell^{p}$. Then the corollary follows from the Five Lemma. ∎
The following definition is introduced by M. Braga, I. Farah, and A. Vignati in Braga et al. [2024].
Definition 4.2 (Braga et al. [2024]).
@@
    If we replace $\mathcal{P}_{Fin}(\prod^{u}_{\mathcal{I}}C^{*}_{u}(X))$ by $\prod^{u}_{\mathcal{I}}C^{*}_{u}(X)$, then we shall say $X$ has operator norm localization, (abbrev. ONL, see Chen et al. [2008]).
We should mention that a large class of spaces satisfies the property $ONL_{\mathcal{P}{Fin}}$, including all metric spaces that admit a coarse embedding into Hilbert space, see [Finn-Sell, 2014, Lemma 35] and [Braga et al., 2024, Theorem 1.3]. Our result extends the class to spaces that are coarsely embeddable into $\ell^{p}$-spaces.
@@
To prove 4.3, we need the following lemma (see also the proof of [Finn-Sell, 2014, Proposition 35]).
@@
Let $X$ be a sparse metric space with bounded geometry. All ghost projections in $C^{*}(X)$ are compact if and only if the canonical inclusion $i:\mathcal{K}\to\mathcal{G}$ induces an isomorphism $i_{*}:K_{0}(\mathcal{K})\to K_{0}(\mathcal{G})$, where $\mathcal{G}=\mathcal{G}(X,X)$ is the ghostly ideal associated with the invariant set $X\subseteq\beta X$.
@@
The $(\Rightarrow)$ part. For any projection $P\in M_{n}(\mathcal{G})$, since the Roe algebra is quasi-stable, there exists a projection $P^{\prime}\in\mathcal{G}$ such that $P]=[P^{\prime}]\in K_{0}(\mathcal{G})$. By assumption, $P^{\prime}$ is compact, thus $P^{\prime}]\in K_{0}(\mathcal{K})$ and $i_{*}([P^{\prime}])=[P]$. Thus $i_{*}$ is surjective. Since $X$ is sparse, the map $i_{*}$ is injective by [Oyono-Oyono and Yu, 2009, Proposition 2.10 & Remark 2.12]. This shows that $i_{*}$ is an isomorphism.
The $(\Leftarrow)$ part. Assume for a contradiction that $P\in C^{*}(X)$ is a non-compact ghost projection. Say $X=\bigsqcup_{n\in\mathbb{N}}X_{n}$ and $N_{n}=\#X_{n}$. Notice that $\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})$ is a subalgebra of $C^{*}(X)$ and $\mathcal{K}(\mathcal{H}_{X})\cap\prod^{u}_{n\in\mathbb{N}}C^{*}(X_{n})=\bigoplus_{n\in\mathbb{N}}M_{N_{n}}(\mathcal{K})$. Moreover, the quotient algebras are canonically isomorphic to each other, i.e.,
@@
Assume for a contradiction that $X$ does not have $ONL_{\mathcal{P}_{Fin}}$. By [Braga et al., 2024, Theorem 1.3], there exists a sparse subspace $X^{\prime}\subseteq X$ such that there exists a non-compact ghost projection in $C^{*}_{u}(X^{\prime})$. Without loss of generality, we may assume that $X$ itself is sparse and $P\in C^{*}_{u}(X)$ is a non-compact ghost projection. Take $p\in\mathcal{K}(\mathcal{H})$ to be a rank one-projection. Then the tensor product $P\otimes p$ defines a non-compact ghost projection in $C^{*}(X)$. By 4.4, the map $i_{*}$ is not an isomorphism on $K_{0}$. This leads to a contradiction to 2.6. ∎